Sol Summary – May 2nd

Crew 261 Sol Summary Report 02-05-2023

Sol: 2

Summary Title: Completion of the crewmembers qualification on EVA procedures on Mars

Author’s name: Aline Decadi, XO

Mission Status: Nominal.

Sol Activity Summary: Today has taken place the EVA-2 with the following crew members: James Burk (Commander) (EVA Leader), Erin Kennedy (Crew Roboticist), Kris Davidson (Crew Journalist). They were accompanied by the crew journalist and the HSO. They prepared our flight suits, backpacks, helmets, and radio. The purpose was training EVA procedures + communication test with crew members and with the hab.
EVA-2 crew members started donning at 8:45 AM and EVA started at 9:15 AM. Our Crew261 Commander pronounced the following emotional beautiful sentences when heading for the first time on Martian soil: “It is with a profound sense of humbleness, duty, and honor that we take these first steps on Mars. We come from Earth, on behalf of all humanity. We come to bring Life to Mars, and Mars to Life”. They drove the Rovers (Spirit and Perseverance) for the first time with their spacesuits to the Marble Ritual, then walked around. We exercised walking around with the spacesuits, making geology/ collecting samples, regular health status, made photographs, and performed communication checks & extensive use of the radio. The communication was confirmed to be good on the field, and the overall safety health checks were nominal. The EVA-2 crew members qualified on EVA procedures today. As a lesson learnt from Sol 1, the EVA total duration was extended from 1h00 to 1h30 to successfully finish on time at 10:22 AM (ETA in the mission request was 10:30 AM).
We spent the afternoon setting up experiments: the GreenHab, the PhotoBioreactor, the Musk observatory first Sun imaging, the file server, the headsets for MarsVR, the robotics Atmosphinder, and cardiovascular monitoring with requested parameters.

We spent the evening having dinner, the commander briefing, and preparing EVA & experiments for tomorrow.

Look Ahead Plan: We will be in Sim tomorrow: soil collection, make a structural test for the Atmosphinder experiment, according to mission request approval.

Anomalies in work: None

Weather: Sunny and pleasant. Very windy day.

Crew Physical Status: As lessons learnt from yesterday, the extended duration for the EVA enables it to function safely and with continuous operational communication. The EVA has been scheduled in the morning as the temperature is more moderate.

EVA: Training EVA to Marble Ritual

Reports to be filed: HSO Report, EVA Request, EVA Report, Operations Report, Astronomy Report, Journalist Report, Pictures of the day.

Support Requested: None.

Operations Report – May 2nd

Crew 261 Operations Report 02-05-2023
SOL: 2
Name of person filing report: Julien Villa-Massone
Non-nominal systems:
– Power system
– Robotic observatory
Notes on non-nominal systems:
Power system: One of the 3 PV strings breaker was found disconnected at 14:30. It is unknown how long this breaker has been disconnected, but a photo showing it connected was taken on April 30 at 13:27 local time. The related MPPT tracker was indicating 0 amps and the fault LED was illuminated. As a consequence, 1/3 of the power was unavailable, and the base was running on 2/3 of the normal PV power. This was reported to Mission Support who promptly reconnected the system. The Crew Engineer will inspect the power system monitoring logs to find out how long this breaker has been disconnected and if a root cause can be determined.
Spirit rover used: Yes
Hours: (before EVA) 220.1
Beginning charge: (Before EVA) 100%
Ending charge: (On return from EVA, before recharging) 87%
Currently charging: Yes
Opportunity rover used: No
Hours: 122.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Curiosity rover used: No
Hours: 232.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Perseverance rover used: Yes
Hours: 262.7
Beginning charge: 100%
Ending charge: 83%
Currently charging: Yes
General notes on rovers: Based on 2 EVAs covering a 1.2km drive each, as noted in yesterday’s report, there are large discrepancies in battery capacity/consumption (in response to yesterday’s report, Dr. Rupert mentioned Percy has older batteries, explaining a lower usable capacity).
Today, we prepared a preliminary usable battery capacity estimate based on an average consumption of 25kWh/100km. This consumption was selected as a preliminary baseline because this is Opportunity’s measured consumption (the best of the 4 rovers), and it is a usual value for sedan/SUV EVs, which comforts the relevance of this value. Preliminary usable battery capacity estimate: Opportunity: ~15kWh (design capacity is 14.9kWh), Curiosity: 5kWh, Spirit: 2.5kWh, Perseverance: 2kWh.
This being preliminary, we will refine these values over time and report.
Summary of Hab operations: Nominal
WATER USE: rate of 33 gallons per day. Cumulative use since Sol 1: 140 Gal (but a suspected faucet leak and the cylindrical shape of the tank contributed to 100 indicated gallons to be lost on Sol 1). At this rate and with today’s water usable volume available, we estimate we will be 2 Sols short, or 68 gallons short.
Water (static tank): 410 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: Nominal
Summary of suits and radios: Nominal. We tested and are now using a satellite positioning and communication system to track EVAs. The primary goal is to improve safety of the crew while on EVA, in two ways. First, a crew member in EVA can now request a rescue (to official rescue services) if a real emergency situation arises by the press of a button. Second, a crew member in EVA can now request support from other crew (in another EVA or at the base) or from mission support at the press of a few buttons. The secondary (but very useful on a daily basis) goal of this system is to enable supporting crew at the Hab to monitor EVA crew in near-real-time (2-min delay at most) on a satellite map, providing relevant context to radio comms. Furthermore, EVA crew can now mark points of interest and send those in near-real-time (<1 min delay) to said map, to add relevant context when discussing a particular location. We will continue to test this system, which will eventually be replaced by the Meshstatic infrastructure.
Summary of GreenHab operations:
WATER USE: 17.8 gallons (Daily water usage for crops: 8,5 gal at 06:40 AM + 6,3 gallons at 07:00 PM. Total = 14,8 gal / Daily water usage for research and/or other purposes: 3 gal.)
Heater: Off
Supplemental light: Off
Harvest: Cucumbers: 1223g, mint: 23g
Summary of ScienceDome operations: Spirulina bioreactor device installation finalized. Culture media preparation completed (8 liters were produced). Culture has been received this evening by cargo. Culture has been partially inserted into the bioreactor (5 liters of media) due to 3 remaining liters of media being still too hot. As a result, 2/3 of the first batch of spirulina culture has been inserted. 1/3 of the first batch is ready for insertion tomorrow. The second batch is standing by as backup. Computer setup is completed and needs to be refined over the next few days to ensure seamless operation (currently, as power loss/restart will not automatically start the experiment).
Biostimulant solution was prepared using spirulina and compost.
Dual split: Off
Summary of RAM operations: (Tools used, work done) Our Crew Robotics Engineer made excellent progress in the assembly of the Atmosphinder robot, using own tools. Assembly is 80% complete. Expecting 100% tomorrow and structural tests.
Summary of any observatory issues: Solar observations have been made, images processed and one stellar image has been produced.
Summary of health and safety issues: Nominal day
Questions, concerns and requests to Mission Support: Thank you for your support throughout the day.

Journalist Report – May 2nd

Crew 261 Journalist Report 01-05-2023

Author: Kris Davidson, Crew Journalist

For early settlers on Mars, there will be a surreal juxtaposition permeating every moment: the staggering responsibility and honor of being among the first humans on Mars, and the unyielding banality of critical daily chores. The extensive list of tasks that fill each Sol is a humbling reminder that even in this far-flung world, life will remain tethered to the mundane.

For Transatlantic Mars 261, these contrasting realities have begun to take hold. Before the EVA on Sol 2, Julien Villa-Massone (Crew Engineer) prepared and checked all the spacesuits — a repetitive task that on Mars could have life or death consequences. EVA 2 was led by James Burk (Commander) with Audrey Derobertmasure (HSO), Erin Kennedy (Crew Robotics Engineer), and Kris Davidson (Crew Journalist) to Marble Ritual for completion of crew training. Safe in the well-maintained spacesuits, the analog astronauts completed the training practice while enjoying an EVA filled with novelty and gentle wonder.

Back at the hab, the crew went about their projects, experiments, and chores. In the morning, and following the EVA, Derobertmasure (HSO) captured biometrics for the cardiovascular aging study. Cécile Renaud (Greenhab Officer) continued work on the biological preparation of spirulina for the photobioreactor experiment, alongside Julien Villa-Massone (Crew Engineer) who attended to the mechanical and programming elements of the device. In the afternoon, Villa-Massone (Crew Engineer) added the EVA Map, a satellite tracking function to the crew website (, for the dual purposes of increasing crew safety while also creating real-time location markers for points of interest. Kennedy (Crew Robotics Engineer) continued programming and assembly of the Atmosphinder wind rover for upcoming EVA testing. Aline Decadi (Executive Officer) worked in the solar observatory to image the sun for education and outreach purposes. Burk (Crew Commander) continued work with the file server and technical set up for the Mars VR mapping work to be completed on an upcoming EVA.

Dinner featured fresh cucumbers, harvested from the greenhab by Renaud (Greenhab Officer). This precious fruit served as an important reminder, relevant to both Earth and Mars: even in the most inhospitable environments, life finds a way to persevere, and the simplest of pleasures can hold profound significance.

GreenHab Report – May 2nd

Crew 261 GreenHab Report 02-05-2023
GreenHab Officer: Cécile Renaud
Environmental control: Door open from 06:40 to 20:15
Average temperatures: (67,6°F at 06:40, 84°F at 10:45, 89°F at 12:20, 91,2°F at 14:30 and 73,8°F at 20:15)
Hours of supplemental light: N/A
Daily water usage for crops: 8,5 gal at 06:40 AM + 6,3 gallons at 07:00 PM. Total = 14,8 gal
Daily water usage for research and/or other purposes: 3 gal.
Water in Blue Tank: 149,2 gallons
Time(s) of watering for crops: 06:40 AM and 20:10 PM
Changes to crops: N/A
Narrative: Nothing to report. Tomatoes will be harvested at SOL 4. Some cucumbers are still growing. Mint/chives are in a very good state, microgreen and tomatoes are less healthy but still in a really good state.
Harvest: (include which crop and mass in grams) Cucumbers: 1223g, mint: 23g
Support/supplies needed: None

EVA Report – May 2nd

Crew 261 EVA Report 02-05-2023

EVA # 2

Author: James Burk, Commander and EVA Lead

Purpose of EVA: EVA and Rover Qualification for crewmembers

Start time: 9:00AM

End time: 10:20 AM

Narrative: We drove the Rovers (Spirit and Perseverance) to the Marble Ritual, then walked around no farther than 200 meters for communication check. We exercised walking around with the spacesuits, regular health status, made photographs, and performed communication checks & extensive use of the radio. The communication was confirmed to be good on the field. Each crew member has used the radio extensively, experienced the spacesuits, the walking, the driving, the team cohesion, and the overall safety health checks were nominal. The EVA-2 crew members qualified on EVA procedures today, some for the second time.

This was James Burk’s first EVA as Commander of a crew at the MDRS, and he dedicated the EVA to Max Boyce, a member of the NorCal chapter who tragically died last year. Max was part of Crew 197 with James in October 2018. Max was a very dedicated Mars Society member and was part of the chapter that works on our EVA suits every year. Max showed James and his 2018 crew how to do EVAs, as well as other duties at MDRS. It was a deeply emotional moment for James.

Destination: Marble Ritual

Coordinates (use UTM WGS 84): 518687 E, 4250789 N

Participants: James Burk (Commander and EVA Lead) (EVA Leader), Erin "RobotZwrrl" Kennedy (Crew Roboticist), Audrey Derobertmasure (HSO), Kris Davidson (Journalist)

Road(s) and routes per MDRS Map: Drive on Cow Dung Road and walk east until Marble Ritual.

Mode of travel: Rovers (Spirit & Perseverance)

Astronomy Report – May 2nd



Crew 261 Astronomy Report 02-05-2023

Astronomy Report

Name: Aline Decadi, XO
Crew: 261
Date: 05/02/2023

Robotic Telescope Requested (choose one) MDRS-14 MDRS-WF MLC => None

Objects to be Imaged this Evening: None
Images submitted with this report: None

Problems Encountered: None


Solar Features Observed: First solar imaging with the Musk Observatory.

The user manual has been operated to perform visual observing the chromosphere, prominences, granules, sunspots, and filaments. The sun observation has been performed first with the zoom eyepiece, then with the camera to captures several thousands of frames and process the images with the following software: AUTOSTAKKERT to stack the images, REGISTAX to add wavelets to bring out fine details, and PHOTOSHOP to merge all in one.
The operation of the dome, telescope and computer were nominal.

Images submitted with this report: Sun 230502 Sunspots.png
Problems Encountered:
-The Sun was moving out the field of view when I was configuring the parameters for capturing a bunch of 1000 frames using the SHARP CAP program, so the adjustment with the red focus knob and the tuning knobs was not easy.
-I used some presets in REGISTAX, but I was not able to adjust efficiently the six wavelets to improve the image processing.
=>These will require more practice.

COMMS open 7-9pm 03May2023

Hello Crew,Mission support is signing in, please submit your daily reports.Please acknowledge.

Weather forecast:
03/05/2023 Night
Temperature – 61F
Humidity – 29%
Cloud Cover – 85%
Wind – S 7 mph
Wind gusts – 27 mph
Probability of Precipitation – 5%

04/05/2023 Morning
Temperature – 56F
Humidity – 42%
Cloud Cover – 35%
Wind – S 6 mph
Wind gusts – 30 mph
Probability of Precipitation – 4%

04/05/2023 Day
Temperature – 71F
Humidity – 14%
Cloud Cover – 19%
Wind – S 19 mph
Wind gusts – 39 mph
Probability of Precipitation – 0%

04/05/2023 Evening
Temperature – 64F
Humidity – 16%
Cloud Cover – 9%
Wind – SW 19 mph
Wind gusts – 36 mph
Probability of Precipitation – 6%

05/05/202307/05/2023 Summary
Daily temperature – 67-74 F
Wind – 15-18 mph
Probability of Precipitation – 1-15%


1) High winds warning tomorrow afternoon. Please plan accordingly.
2) FYI, an earth vehicle is coming tomorrow morning to empty the septic tank.
Sergii Iakymov – Assistant Director Mars Desert Research Station


Copyright © The Mars Society. All rights reserved. | Main Site